# Distributed Differentially Private Data Analytics via Secure Sketching

### v1

@@ 3. Privacy Guarantees in the LTM
We consider a model in which the trusted component can perform any public linear transformation of the inputs. The model includes three algorithms: (1) $R:\mathcal{X}\rightarrow\mathcal{Y}$ is a randomized encoder that takes a single user’s data and outputs a randomized message, (2) $T:\mathcal{Y^{*}}\rightarrow\mathcal{Y^{*}}$ is the idealized trusted component that performs a public linear transformation of its inputs, and (3) $A:\mathcal{Y^{*}}\rightarrow\mathcal{Z}$ is an analysis function that takes the results messages and estimates some function from these messages.
Note that standard definitions of differential privacy in the shuffle model only require $T(R(x_{1}),...,R(x_{n}))$ to be differentially private. This implicitly assumes that all clients are honest and do not collude with the adversary, in particular they are assumed not to leak the output of their randomizers publicly. This implies that existing definitions of differential privacy in the shuffle model could be satisfied even by (artificial) mechanisms in which a single client adds the whole noise while the others do not randomize their messages at all. This is clearly a very weak privacy guarantee: in a setting in which a large number of clients participate in a differentially private data analysis, it is unrealistic to assume that the adversary does not control even a single client. Luckily, to the best of our knowledge, no proposed mechanisms in the literature suffer from these vulnerabilities, still this counterexample shows that the existing definition is not robust enough, and we therefore formalize a notion of differential privacy where we explicitly tolerate that a bounded number of clients might collude with the adversary. Similar observations were also made by (TWM+, 23).
@@
This definition requires any possible subset of honest parties to jointly add enough randomness in $R$ to preserve the privacy guarantee, thus guaranteeing privacy even against a semi-honest adversary that is able to learn the outputs of an arbitrary subset of clients.
@@
(Ste, 20) introduce the multi-central model, which exactly defines this split-trust model that instantiates the trusted computation model above. They allow a semi-honest adversary that honestly follows the protocol to corrupt up to all but one servers and all but one client. We also operate in the same semi-honest setting, but define parameters $t$ and $t^{\prime}$ for the number of tolerated server and client corruptions.
The view $\mathtt{View}^{R,\Pi,A}_{C_{cor},S_{cor}}(x)$ of the adversary consists of all the information available to the corrupted clients $C_{cor}$ and servers $S_{cor}$ plus the final output of the honest, uncorrupted servers. This view excludes only internal information of the trustworthy servers and clients. The resulting protocol is $(\varepsilon,\delta)$-differentially private in the multi-central model if adversary’s view is $(\varepsilon,\delta)$-indistinguishable from the output on a neighboring dataset.
Figure 1 further illustrates what is contained in the adversary’s view. All clients randomize their input and send a single message to each server, who locally perform a linear transformation and release the result to some analyzer. If one server and one client are corrupted, the adversary’s view contains all incoming and outgoing messages from that server and that client, or the incoming and outgoing edges from the client, server, and analyzer marked in red in the figure.
@@
Let $\Pi$ be a $k$-party MPC protocol that computes $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with perfect security, tolerating $t$ corruptions. A tuple of algorithms $P=(R,\Pi,A)$ is $(\varepsilon,\delta)$-differentially private if for all coalitions $S_{cor}$ of up to $t<k$ corrupt servers and for all coalitions $C_{cor}$ of $t^{\prime}<n$ corrupt clients and all neighboring datasets $x$ and $x^{\prime}$:
$\Pr[\mathtt{View}^{R,\Pi,A}_{S_{cor},C_{cor}}(x)\in S]\leq e^{\varepsilon}\cdot\Pr[\mathtt{View}^{R,\Pi,A}_{S_{cor},C_{cor}}(x^{\prime})\in S]+\delta$
@@
With this definition, we can give our first result showing that differential privacy is retained given a bounded number of corruptions. The proof can be found in Appendix B.
Lemma 3.3.
If $P=(R,T,A)$ is $(\varepsilon,\delta)$-differentially private with $t^{\prime}$ corrupt clients $C_{cor}$, and if $\Pi$ is a perfectly secure $k$-party MPC protocol that computes $T$ correctly while tolerating $t$ corrupt servers $S_{cor}$, then $P=(R,\Pi,A)$ is $(\varepsilon,\delta)$-differentially private as long as there are less than $t^{\prime}$ corrupt clients and $t$ corrupt servers.
3.1. Frequency Moments
To illustrate the possibilities inherent to the LTM, we consider a simple application by way of estimating frequency moments. For simplicity, assume in this section that all clients are trustworthy, i.e. $t^{\prime}=0$, but the arguments can be straightforwardly extended to deal with arbitrary values of $t^{\prime}$. Here, each client is given a number $[-\Delta,\Delta]$ and our goal is to estimate $F_{k}=\sum_{i=1}^{n}|x_{i}|^{k}$. In the local model, even if all entries are either 00 or $1$, it is not possible to estimate $F_{1}$ without incurring an additive error of the order $\sqrt{n}$, that is the estimated value $\tilde{F_{1}}=F_{1}\pm O(\sqrt{n})$ for any privacy preserving local mechanism (BS, 15). In contrast, simply computing $F_{k}$ and adding an appropriate amount of noise [^1] Using the Gaussian mechanism, $\sigma^{2}=\rho\cdot\frac{\Delta^{k}}{\varepsilon^{2}}\log\delta^{-1}$ for some absolute constant $\rho$ is sufficient, though other mechanism can yield even better bounds. yields an additive error that only depends on $\Delta$, $\varepsilon$ and $\delta$.
We now consider the LTM. Let $\mathbf{x}^{k}$ denote the vector of client entries with $\mathbf{x}^{k}_{i}=|x_{i}^{k}|$. We observe that $F_{k}=1^{T}\mathbf{x}^{k}$. Suppose every client $i$ samples a Gaussian random variable $g_{i}$ with zero mean and variance $\sigma^{2}\geq\frac{2\Delta^{k}\ln(1.25/\delta)}{n\cdot\varepsilon^{2}}$ and adds it to $|x_{i}^{k}|$. Then we have $\tilde{F_{k}}=1^{T}(\mathbf{x}^{k}+g)=F_{k}+\sum_{i=1}^{n}g_{i}$. We now observe that $\sum_{i=1}^{n}g_{i}$ is Gaussian distributed with zero mean and variance $n\cdot\sigma^{2}\geq\frac{2\Delta^{k}\ln(1.25/\delta)}{\varepsilon^{2}}$. Thus, assuming a trusted computation of $1^{T}g$, $\tilde{F_{k}}$ is a differentially private estimate of $F_{k}$ with error $\frac{2\Delta^{k}\ln(1.25/\delta)}{\varepsilon^{2}}$.
A key property used in this example (and throughout the paper), is that sums of Gaussian random variables are Gaussian distributed. A similar property also holds for Cauchy random variables, or random variables drawn from any stable distribution. Thus, differential privacy protocols based on the Gaussian or the Cauchy mechanism (NRS, 07) are natural baseline candidates for the LTM. Nevertheless, it is also possible to add noise following a distribution where the sum of random variables can be controlled. For example, the sum of negative binomial random variables follows a geometric distribution, which likewise can be used in differential private mechanisms (GKMP, 20). Exploring these options in the context of the LTM is a interesting open problem.
3.2. Privacy Preserving Mechanism for Oblivious Sparse Johnson Lindenstrauss Transforms
We instantiate the trusted computation model for differential privacy using MPC, where the function $T$ performs a matrix multiplication and outputs a sketch of the input data. More specifically, we define the tuple of algorithms $(R_{\sigma},T_{\mathbf{S}},A)$. $R_{\sigma}(\mathbf{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{ds}$ sets the input value to $0^{d}$ if $n<8m\ln(dm/\delta)+t^{\prime}$. Otherwise, for all $i\in[s]$, sample $\mathbf{g}_{i}\leftarrow\mathcal{N}(0,\sigma^{2})^{d}$ with variance $\sigma^{2}$, and add these sample vectors to the input vector, resulting in $s$ vectors $\mathbf{x}_{i}=\mathbf{x}+\mathbf{g}_{i}$. Each client then secret shares the resulting values using a linear secret sharing scheme and sends one share to each server. $T_{\mathbf{S}}:\mathbb{R}^{n\times ds}\rightarrow\mathbb{R}^{m\times d}$ has as a parameter sketching matrix $\mathbf{S}\leftarrow\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ and takes as input matrices $\mathbf{A}_{i}\in\mathbb{R}^{n\times d}$ for $i\in[s]$. Notice that $\mathbf{S}$ satisfying Definition 2.1 can be decomposed such that $\mathbf{S}=\frac{1}{\sqrt{s}}\sum_{i}^{s}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ is a matrix with only one non-zero entry per column. $T_{\mathbf{S}}$ outputs the sum of matrix products $\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{A}_{i}$. $A$ takes a matrix in $\mathbb{R}^{m\times d}$, as well as any public values, and outputs some data analysis.
@@ 3.2. Privacy Preserving Mechanism for Oblivious Sparse Johnson Lindenstrauss Transforms
2:  Parameters: noise variance $\sigma^{2}$, public sketching matrix $\mathbf{S}\in\mathbb{R}^{m\times n}$
3:  Each client $i\in[n]$ then locally computes $R_{\sigma}(\mathbf{x}_{i})$ on their input $\mathbf{x}_{i}$.
@@
4:  Each client secret shares the resulting value using a linear secret sharing scheme, and sends one share to each server.
@@
5:  The servers jointly compute function $T_{\mathbf{S}}$ on the resulting matrix, which results from concatenating all $n$ vector secret shares.
@@
6:  The resulting matrix product is be published and then taken as input to any analysis function $A$.
7:  Output: Output $A(T_{\mathbf{S}}(R_{\sigma}(\mathbf{x}_{1},\dots,\mathbf{x}_{n}))$
@@
Algorithm 1 describes how clients and servers jointly compute the linear transformation, allowing the transformed data to be published and used for analysis. We now evaluate the variance necessary to guarantee differential privacy in this model. We find that for sufficiently large $n$, we can apply the stability of Gaussian distributions to divide by the number $(n-t^{\prime})/2m$ of honest clients expected to contribute to any entry of a column in the resulting noisy matrix product. The concentration bound, formulated in Lemma B.1 and used to bound this number of honest clients, contributes to the necessary value of $\delta$. Composition theorems can be applied on the columns of the resulting matrix product to yield the final differential privacy guarantee. We will later see that adding noise with this variance to data results in only a small multiplicative error in the resulting error for regularized linear regression and low rank approximation.
We are now ready to state our privacy guarantees in the LTM based on the Gaussian mechanism.
@@
Let $\varepsilon\geq 0$, $\delta\in(0,1)$, $t^{\prime}<n$, $m\in[n]$, $s\in[n]$ and
$\sigma^{2}=\frac{4s^{2}\eta^{2}\ln(1.25s/(\delta/d-m\exp(\frac{-(n-s-t^{\prime})}{8m})))m^{2}d^{2}}{\varepsilon^{2}(n-s-t^{\prime})}$
@@
Let $\mathbf{S}\sim\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ with $t$ non-zero entries per column. Then as long as input values are bounded above by $\eta$, tuple of algorithms $(R_{\sigma},T_{\mathbf{S}},A)$ is $(\varepsilon,\delta)$-differentially private in the trusted computation model for differential privacy with $t^{\prime}$ corrupt clients $C^{\prime}$.
The proofs can be found in Appendix B. Corollary 3.5 below follows directly from Theorem 3.4 and Lemma 3.3.
Corollary 3.5.
Let $(R_{\sigma},T_{\mathbf{S}},A)$ be the tuple of algorithms above. If $T_{\mathbf{S}}$ is computed correctly using an MPC protocol $\Pi_{\mathbf{S}}$ with perfect security tolerating $t$ semi-honest corruptions $S_{cor}$, then the tuple of algorithms $(R_{\sigma},\Pi_{\mathbf{S}},A)$ is $(\varepsilon,\delta)$-differentially private with $t^{\prime}$ corrupt clients $C_{cor}$.
3.3. Cryptographic Assumptions and Relations to the Shuffle Model
Crucially, the execution of the linear transformation can be easily and securely distributed using simple cryptographic techniques for secure multiparty computation (MPC) such as linear secret-sharing (LSS). The guarantee offered by MPC is that no adversary corrupting all but one of the servers can learn more than the output of the computation performed. (DKM+, 06) is the first work to consider the combination of differential privacy and MPC; their focus is distributed noise generation, which requires interaction between servers. (CY, 23) provide lower bounds for non-interactive multi-server mechanisms. (DKN+, 23) provide an interactive MPC protocol to compute selection for distributed trust models of differential privacy. The linear transforms can be non-interactively computed locally by servers and are also known to all participating parties, that is, once initiated, the output of the protocol is fully deterministic. This allows to implement the LTM with a set of central servers under the minimal assumption that at least one honest server will not collude with all the others, without any requirement to fully trust any particular server. Using only simple LSS has several important consequences for both the efficiency and the security of the overall system: each client only needs to send a single message to each server, and servers do not need to communicate with each other, but only send a single message to the data analyst. Thus, the total workload is essentially that of the central model times the number of servers used. Finally, as LSS can be instantiated with information-theoretic security, the security of our system does not depend on any unproven mathematical assumption and immediately offers security even against powerful quantum adversaries. We describe the type of secret sharing scheme that can be used to securely compute linear transformations using MPC in Appendix E.
Secure aggregation and the shuffle model are two instantiations of intermediate trust models for differential privacy. Secure aggregation (GX, 17; BIK+, 17; MPBB, 19; AG, 21; TWM+, 23) is a special case of the LTM, where the linear transformation applied is a sum, which is useful in federated learning. The shuffle model (BEM+, 17; EFM+, 19; CSU+, 19) is motivated in practice by the existence of cryptographic protocols (e.g., mixnets) which allow to securely distribute a shuffle among a set of servers. However, implementing shuffles using cryptography is significantly more cumbersome than implementing linear transformations. In particular mixnets require: 1. servers to talk to each other over a chain, thus increasing the latency of the system (while in the LTM all servers can perform the computation in parallel and without talking to each other); 2. the use of computationally intensive public-key cryptography (while LSS only requires performing linear operations); 3. to use computational assumptions, (while LSS provides unconditional security). Finally, shuffles/mixnets are inherently randomized algorithms, while linear transformations are deterministic functions and as such are simpler to implement securely. [^2] The linear transformations we use are random, but this randomness is not private, so we don’t need to trust the servers not to leak it. Instead, linear transformations are sampled before secure computation.
4. Numerical Linear Algebra in LTM
In this section we give the utility guarantees of our mechanism. We start with giving generic distortion bounds relating the spectrum of the noisy, but private matrix to the the spectrum of the original non-private matrix. Parameterizations of this mechanism depend on the underlying class of subspace embeddings. In this utility analysis, we limit outselves to OSNAPs with $m$ and $s$ chosen according to (Coh, 16); other trade-offs are possible and can be found in Appendix A. Applications to specific problems such as regression and low rank approximation are given towards the end of the section.
Here we present the formal utility statements for ridge regression and low rank approximation. All bounds have a dependency on $\sigma^{2}n$. This term is in $\operatorname{poly}(d\varepsilon^{-1}\log(1/\delta))$ by our parametrization of the Gaussian mechanism. All proofs can be found in the Appendix C.
We will use the following spectral bounds for both linear regression as well as low rank approximation. We believe that there may be further applications and that the bounds themselves are therefore of independent interest.

### v2

@@ 4. Privacy Guarantees in the LTM
We consider a model in which the trusted component can perform any public linear transformation of the inputs. The model includes three algorithms: (1) $R:\mathcal{X}\rightarrow\mathcal{Y}$ is a randomized encoder that takes a single user’s data and outputs a randomized message, (2) $T:\mathcal{Y^{\prime}}\rightarrow\mathcal{Y^{*}}$ is the idealized trusted component that performs a public linear transformation of its inputs, and (3) $A:\mathcal{Y^{*}}\rightarrow\mathcal{Z}$ is an analysis function that takes the results messages and estimates some function from these messages.
Note that standard definitions of differential privacy in the shuffle model only require $T(R(x_{1}),...,R(x_{n}))$ to be differentially private. This implicitly assumes that all clients are honest and do not collude with the adversary, in particular they are assumed not to leak the output of their randomizers publicly. This implies that existing definitions of differential privacy in the shuffle model could be satisfied even by (artificial) mechanisms in which a single client adds the whole noise while the others do not randomize their messages at all. This is clearly a very weak privacy guarantee: in a setting in which a large number of clients participate in a differentially private data analysis, it is unrealistic to assume that the adversary does not control even a single client. Luckily, to the best of our knowledge, no proposed mechanisms in the literature suffer from these vulnerabilities, still this counterexample shows that the existing definition is not robust enough, and we therefore formalize a notion of differential privacy where we explicitly tolerate that a bounded number of clients might collude with the adversary. Similar observations were also made by (Talwar et al., 2023).
@@
(Steinke, 2020) introduce the multi-central model, which exactly defines this split-trust model that instantiates the trusted computation model above. They allow a semi-honest adversary that honestly follows the protocol to corrupt up to all but one servers and all but one client. We also operate in the same semi-honest setting, but define parameters $t$ and $t^{\prime}$ for the number of tolerated server and client corruptions.
The view $\mathtt{View}^{R,\Pi,A}_{C_{cor},S_{cor}}(x)$ of the adversary consists of all the information available to the corrupted clients $C_{cor}$ and servers $S_{cor}$ plus the final output of the honest, uncorrupted servers. This view excludes only internal information of the trustworthy servers and clients. The resulting protocol is $(\varepsilon,\delta)$-differentially private in the multi-central model if the adversary’s view is $(\varepsilon,\delta)$-indistinguishable from the output on a neighboring dataset.
Figure 1 further illustrates what is contained in the adversary’s view. If one server and one client are corrupted, the adversary’s view contains all incoming and outgoing messages from that server and that client marked in red in the figure.
@@
Let $\Pi$ be a $k$-party MPC protocol that computes $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with perfect security, tolerating $t$ corruptions. A tuple of algorithms $P=(R,\Pi,A)$ is $(\varepsilon,\delta)$-differentially private if for all coalitions $S_{cor}$ of up to $t<k$ corrupt servers and for all possible $S$, all coalitions $C_{cor}$ of $t^{\prime}<n$ corrupt clients and all neighboring datasets $x$ and $x^{\prime}$:
$\Pr[\mathtt{View}$ (${}^{R,\Pi,A}_{S_{cor},C_{cor}}(x)\in S]$
$\leq e^{\varepsilon}\cdot\Pr[\mathtt{View}^{R,\Pi,A}_{S_{cor},C_{cor}}(x^{\prime})\in S]+\delta$
@@
Algorithm 1 describes how clients and servers jointly compute the linear transformation, allowing the transformed data to be published and used for analysis.
@@
2:  Each client $i\in[n]$ locally computes $R(\mathbf{x}_{i})$ on their input $\mathbf{x}_{i}$.
@@
3:  Each client secret shares the resulting value using a linear secret sharing scheme, and sends one share to each server.
@@
4:  The servers jointly compute function $T$ on the matrix, which results from concatenating all $n$ vector secret shares.
@@
5:  The resulting matrix product is published and then taken as input to any analysis function $A$.
6:  Output: Output $A(T(R(\mathbf{x}_{1}),\dots,R(\mathbf{x}_{n}))$
With this definition, we can give our first result showing that differential privacy is retained given a bounded number of corruptions.
Lemma 4.3.
If $P=(R,T,A)$ is $(\varepsilon,\delta)$-differentially private with $t^{\prime}$ corrupt clients, and if $\Pi$ is a perfectly secure $k$-party MPC protocol that computes $T$ correctly while tolerating $t$ corrupt servers, then $P=(R,\Pi,A)$ is $(\varepsilon,\delta)$-differentially private.
Proof Sketch..
This proof can be seen as analogous to the security proof of Theorem 4 in (Talwar et al., 2023). Recall that each server learns nothing more from a perfectly secure multi-party computation protocol other than what is implied by their own input and the output of the function being evaluated. Then there is a simulator $\mathtt{Sim}^{R,\Pi,A}_{S_{cor},S_{cor}=c}(x)$ whose output is $(\varepsilon,\delta)$ indistinguishable from the adversary’s view $\mathtt{View}^{R,\Pi,A}_{S_{cor},C_{cor}}(x)$. Note that the secret shares that are part of the view of the corrupted servers in the protocol $\Pi$ are independent of the input values of the corrupted clients, and all can be easily simulated by random sampling.
A bit more formally, let $x$ be an input dataset, where $x_{i}$ is the input value contributed by client $i$ and $\tilde{x}_{i}$ is the vector of input shares for server $i$ that serves as input to $\Pi$. There is a mapping from $x\in\mathcal{X}^{n}$ to $\tilde{x}\in\tilde{\mathcal{X}}^{k}$. Denote $X=\{x_{i}\}_{i\in C_{cor}}$ and $X^{\prime}=\{x^{\prime}_{i}\}_{i\in C_{cor}}$.Then for any two datasets $x$, $x^{\prime}$ that differ in one entry $x_{i}$ with $i\notin C_{cor}$:
$\Pr[\mathtt{View}^{R,\Pi,A}_{S_{cor},C_{cor}}(x)\in S]$
$\leq\Pr[\mathtt{Sim}^{R,\Pi,A}_{S_{cor},C_{cor}}(f(R(x_{1}),\dots,R(x_{n}),X)\in S]$
$\leq e^{\varepsilon}\Pr[\mathtt{Sim}^{R,\Pi,A}_{S_{cor},C_{cor}}(f(R(x^{\prime}_{1}),\dots,R(x^{\prime}_{n}),X^{\prime})\in S]+\delta$
4.1. Frequency Moments
To illustrate the possibilities inherent to the LTM, we consider a simple application by way of estimating frequency moments. For simplicity, assume in this section that all clients are trustworthy, i.e. $t^{\prime}=0$, but the arguments can be straightforwardly extended to deal with arbitrary values of $t^{\prime}$. Here, each client is given a number $[-\Delta,\Delta]$ and our goal is to estimate $F_{k}=\sum_{i=1}^{n}|x_{i}|^{k}$. In the local model, even if all entries are either 00 or $1$, it is not possible to estimate $F_{1}$ without incurring an additive error of the order $\sqrt{n}$, that is the estimated value $\tilde{F_{1}}=F_{1}\pm O(\sqrt{n})$ for any privacy preserving local mechanism (Bassily and Smith, 2015). In contrast, simply computing $F_{k}$ and adding an appropriate amount of noise yields an additive error that only depends on $\Delta$, $\varepsilon$ and $\delta$.
We now consider the LTM. Let $\mathbf{x}^{k}$ denote the vector of client entries with $\mathbf{x}^{k}_{i}=|x_{i}^{k}|$. We observe that $F_{k}=f_{k}(\mathbf{x})=1^{T}\mathbf{x}^{k}$. Suppose that every client $i$ samples $g_{i}$ from a distribution, and we denote $\tilde{\mathbf{x}}$ the vector of noisy client entries with $\tilde{\mathbf{x}}^{k}_{i}=|x_{i}^{k}|+g_{i}$. To guarantee differential privacy, we can use a differentially private additive noise mechanism $1^{T}(\tilde{\mathbf{x}}^{k})=f_{k}(\mathbf{x})+G$, where the noise distribution is infinitely divisible and $G=\sum_{i\in[n]}g_{i}$.
Many different infinitely divisible distributions can be used at this stage. One example is the geometric distribution, since the sum of negative binomial random variables follows a geometric distribution, which can be used in differentially private mechanisms (Ghazi et al., 2020). The Laplace distribution, which is commonly used for $\varepsilon$-DP mechanisms, is also infinitely divisible, since a zero-centered Laplace distribution can be formulated as a sum of differences between Gamma distributions. Other possible options are stable distributions, a subset of infinitely divisible distributions for which each $g_{i}$ and $G$ follow the same distribution, such as Gaussian or Cauchy.
Since Gaussians are popular in differentially private mechanisms, we will describe the parameter choices for this distribution. Let $g_{i}$ be Gaussian distributed with zero mean and variance $\sigma^{2}\geq\frac{2\Delta^{k}\ln(1.25/\delta)}{n\cdot\varepsilon^{2}}$. Then we have $\tilde{F_{k}}=1^{T}(\mathbf{x}^{k}+g)=F_{k}+\sum_{i=1}^{n}g_{i}$. We now observe that $\sum_{i=1}^{n}g_{i}$ is Gaussian distributed with zero mean and variance $n\cdot\sigma^{2}\geq\frac{2\Delta^{k}\ln(1.25/\delta)}{\varepsilon^{2}}$. Thus, assuming a trusted computation of $1^{T}g$, $\tilde{F_{k}}$ is a differentially private estimate of $F_{k}$ with error $\frac{2\Delta^{k}\ln(1.25/\delta)}{\varepsilon^{2}}$.
We also describe the analogous parameter choices for the Laplace mechanism. Let $g_{i}$ be sample from the distribution defined as the difference between two Gamma distributions $\Gamma(\frac{1}{n},\frac{2\Delta^{k}}{\varepsilon})$. Again, we have $\tilde{F_{k}}=1^{T}(\mathbf{x}^{k}+g)=F_{k}+\sum_{i=1}^{n}g_{i}$. We now observe that $\sum_{i=1}^{n}g_{i}$ is Laplace distributed, centered at 0 and with scale $\frac{2\Delta^{k}}{\varepsilon}$. Thus, assuming a trusted computation of $1^{T}g$, $\tilde{F_{k}}$ is a differentially private estimate of $F_{k}$ with error $\frac{2\Delta^{k}}{\varepsilon}$.
4.2. Privacy Preserving Mechanism for Oblivious Johnson Lindenstrauss Transforms
We instantiate the trusted computation model for differential privacy for specific choices for the functions in tuple $(R,T,A)$. We will show how to guarantee differential privacy for the tuples $(R_{D}^{\text{dense}},T_{\mathbf{S}}^{\text{dense}},A)$ and $(R_{D},T_{\mathbf{S}},A)$, which we will refer to as the dense sketching mechanism and the sparse sketching mechanism, respectively.
Randomizer function $R_{D}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{ds}$. The randomizer takes as input a $d$-dimensional row vector $\mathbf{x}$, copies it $s$ times and concatenates the result to form $\mathbf{x}^{s}$. Then noise is sampled independently from distribution $D$ and added to each of the $ds$ entries in the resulting vector.
(1) $R^{\text{dense}}_{D}(\mathbf{x})=\mathbf{x}^{s}+\mathbf{g},\mathbf{g}\sim D^{ds}$
We define a second alternative to this randomizer function, which instead outputs $0^{ds}$ if $n<8m\ln(dm/\delta)+t^{\prime}$:
(2) $R_{D}(\mathbf{x})=\begin{cases}0^{ds}&n<8m\ln(dm/\delta)+t^{\prime}\\ R^{\text{dense}}_{D}(\mathbf{x})&else\end{cases}$
Transformation function $T_{\mathbf{S}}:\mathbb{R}^{n\times ds}\rightarrow\mathbb{R}^{m\times d}$. This function takes as a parameter a matrix $\mathbf{S}\in\{-1,0,1\}^{m\times n}$ and takes as input a matrix $\mathbf{A}\in\mathbb{R}^{n\times ds}$, which corresponds to the vertical concatenation or stacking of $n$ outputs from a randomizer function $R$. We parse $\mathbf{A}$ as the horizontal concatenation of matrices $\mathbf{A}_{i}\in\mathbb{R}^{n\times d}$ for $i\in[s]$. Let $s$ be the largest number of non-zero entries in a column of $\mathbf{S}$. Then $\mathbf{S}$ can be decomposed such that $\mathbf{S}=\frac{1}{\sqrt{s}}\sum_{i}^{s}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ is a matrix with only one non-zero entry per column, and the assignment of non-zero entries to indeces $i$ is done uniformly at random. $T_{\mathbf{S}}$ outputs the sum of matrix products $\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{A}_{i}$.
(3) $T_{\mathbf{S}}(\mathbf{A})=\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{A}_{i}$
We also define $T_{\mathbf{S}}^{\text{dense}}:\mathbb{R}^{n\times dm}\rightarrow\mathbb{R}^{m\times d}$, where $\mathbf{S}$ is a dense matrix rather than a sparse one, so we set $s=m$. Then $\mathbf{S}$ can be decomposed such that $\mathbf{S}=\frac{1}{\sqrt{m}}\sum_{i}^{m}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ is a matrix with only one non-zero entry per column and exactly $n/m$ non-zero entries per row.
Analysis function $A$. $A$ takes a matrix in $\mathbb{R}^{m\times d}$, as well as any public values, and outputs some data analysis. Examples include low rank approximation and linear regression.
@@ 4.2. Privacy Preserving Mechanism for Oblivious Johnson Lindenstrauss Transforms
We now evaluate the noise distribution necessary to guarantee differential privacy in this model. We find that for sufficiently large $n$, we can apply the infinite divisibility of distributions to divide the noise into a number of pieces corresponding the number $(n-t^{\prime})/2m$ of honest clients expected to contribute to any entry of a column in the resulting noisy matrix product. The concentration bound, formulated in Lemma 4.5 and used to bound this number of honest clients, contributes to the necessary value of $\delta$. Composition theorems can be applied on the columns of the resulting matrix product to yield the final differential privacy guarantee. We will later see that adding suitable noise to data results in only a small error in the resulting error for regularized linear regression and low rank approximation.
We are now ready to state our privacy guarantees in the LTM.
@@
Let $\varepsilon\geq 0$, $\delta\in(0,1)$, $m\in[n]$, $t^{\prime}<n$, $s\in[m]$. Let $\mathbf{S}\sim\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$ with $s$ non-zero entries per column. Also let noise $g_{i}$ be sampled from a symmetric distribution $D$ such that $\sum_{i\in[(n-s-t^{\prime})/(2m)]}g_{i}=G$. $G$ follows a distribution $D^{\prime}$ that defines an additive noise mechanism that satisfies $(\varepsilon/(sd),\delta/(sd)-m\exp(-(n-s-t^{\prime})/8m)))$-DP for function $f(\mathbf{A})=(\mathbf{S}\mathbf{A})_{j}+\mathbf{n}$ for $\mathbf{n}\sim D^{\prime m}$ on input $\mathbf{A}\in\mathbb{R}^{n\times d}$, where the subscript refers to one column of $\mathbf{S}\mathbf{A}$. Then as long as input values are bounded above by $\eta$, the sparse sketching mechanism $(R_{D},T_{\mathbf{S}},A)$, as defined by equations 2 and 3 and for any $A$, is $(\varepsilon,\delta)$-differentially private in the trusted computation model for differential privacy with $t^{\prime}$ corrupt clients.
Proof.
We begin by proving a helper lemma below.